# Literature Review: Cyber Security Monitoring in Maritime

### v2

@@ Abstract
In recent years, many cyber incidents have occurred in the maritime sector, targeting the information technology (IT) and operational technology (OT) infrastructure. Although several literature review papers have been published in the maritime field, none of the previous studies has focused on cyber security monitoring, which aims at timely detection of cyber attacks with automated methods. The current article addresses this research gap and surveys the methods, algorithms, tools and architectures used for cyber security monitoring in the maritime sector. For the survey, a systematic literature review of cyber security monitoring studies is conducted. The first contribution of this article is the bibliometric analysis of related literature and the identification of the main research themes in previous works. For that purpose, our article presents a taxonomy for existing studies which highlights the main properties of maritime cyber security monitoring research. The second contribution of this article is an in-depth analysis of previous works and the identification of research gaps and limitations in existing literature. Based on our findings, we outline future research directions for cyber security monitoring in the maritime field.
@@
cyber security, cyber security monitoring, intrusion detection, literature review, maritime, monitoring
@@ I Introduction
One of the key approaches for countering cyber attacks is cyber security monitoring (or security monitoring for the sake of brevity), which aims at detecting cyber attacks with automated methods in a timely fashion, allowing to mitigate the attacks in a manual, semi-automated, or fully automated way. For example, network and host-based Intrusion Detection Systems (IDSs) are widely acknowledged security monitoring technologies for real-time surveillance of computer networks and individual hosts to detect cyber attacks. Also, most security-aware organisations operate a Security Operations Centre (SOC) for collecting data from security monitoring tools to a central Security Information and Event Management (SIEM) system where it is processed and then presented to human security analysts for further action [16]. In addition to existing industrial solutions, many experimental approaches like Machine Learning (ML) based algorithms have been proposed for cyber security monitoring in recent academic literature (see a domain overview paper by Apruzzese et al. [17] for more details).
@@ II Related work
Clavijo Mesa et al. [19] conducted an SLR on cyber attacks in maritime supply chain networks (MSCN), providing recommendations on the prevention and mitigation practices. The paper introduced a comprehensive taxonomy for cyber attacks which categorised them based on a number of factors, including the hacker group, attack type, affected systems, geopolitical impact, etc. According to the study, the key cyber security threats to MCSN include malware, Distributed Denial of Service (DDoS) and Denial of Service (DoS) attacks. For prevention and mitigation, the paper proposed a number of technical and policy-level measures (e.g., the use of multi-factor authentication, standardisation of cyber security practices, etc.), and practices for increasing cyber security awareness (e.g., development of training courses about attack prevention). The study briefly mentioned some security monitoring technologies without a deeper analysis.
@@
Yu et al. [22] reviewed the papers on cyber security in the maritime domain, describing common trends in relevant literature. As commonly occurring research themes, the authors identified the impact assessment for maritime cyber threats, recognition of maritime cyber threats (e.g., through vulnerability management tools or through interviews with human personnel), and categorisation and mitigation of maritime cyber risks. The study also presented recommendations for risk mitigation, and discussed gaps in existing literature and avenues for further research. As a supplementary material to their paper, the authors published a list of 21 maritime cyber attacks that have received media coverage and public attention, providing a short description of each cyber attack. Although the study briefly mentioned couple of security monitoring approaches for cyber risk mitigation (e.g., the use of ML methods for AIS anomaly detection), these approaches were not elaborated further.
Afenyo and Caesar [7] focused on maritime cyber security threats in their overview paper, analysing relevant literature in the field. According to the authors, existing papers do not provide enough data on past cyber security incidents and this complicates research on modelling these attacks. In addition, the authors stated that the research on maritime cyber attack data-sharing methods has received little attention so far. Furthermore, it was stated that existing studies have not focused on assessing the financial impact of cyber attacks. The authors also noted that the current educational system does not prepare maritime professionals well enough to handle cyber threats. Finally, the authors pointed out that existing governmental and international policies do not regulate maritime cyber security well enough. For example, regulations should facilitate timely reporting of cyber security incidents and cooperation between maritime companies to increase knowledge sharing on the nature and mitigation of cyber attacks.
Erbas et al. [20] performed an SLR that critically examined and compared existing threat modelling and risk assessment methods in ship cyber security, developing a taxonomy for them. The authors analysed 25 scientific papers to understand the evolving landscape of cyber security practices for manned and autonomous ships. Significant inconsistencies were observed in current approaches and key challenges were identified, underscoring the urgent need for standardised threat modelling and risk assessment frameworks that consider also the special requirements of autonomous ships. Furthermore, the authors highlighted disparities, as well as inadequacy in tool support for the reviewed threat modelling and risk assessment methods. The need for more systematic validation of expert knowledge was also deemed critical. The topic reviewed by Erbas et al. is complementary to security monitoring, as threat modelling and risk assessment can provide direct inputs for enhanced monitoring solutions.
@@
A bibliometric review by Chaal et al. [23] analysed the research literature on the risk, reliability, and safety of autonomous ships. Whereas cyber security was not the primary focus of this paper, it was identified as an important part of the ship’s safety. Also, the study indicated that the cyber security of autonomous vessels has started to receive more attention from researchers recently. However, the paper did not include the treatment of cyber security monitoring topics.
Another study by Ben Farah et al. reviewed cyber incidents in the maritime domain and provided a concise bibliographic analysis [4]. The authors provided an extensive elaboration of maritime sub-systems located in vessels and ports, in addition to a review of the potential cyber attacks targeting these systems. Similarly to previous studies described in this section, the authors did not focus on cyber security monitoring issues.
@@ III-B Survey Method
During the development of the review the PRISMA protocol [25, 26] was followed. According to the protocol, an SLR should have clearly defined objectives or research questions (see Section III-A), eligibility criteria for relevant literature (discussed in Section III-B1), a search strategy for finding relevant studies and selection process (discussed Section III-B2), and data extraction process (discussed in Section III-B3).
@@ III-B2 Search Strategy
The automated search from all libraries resulted in over 1000 papers. After the library search, the four co-authors processed one library each in the form of independent work, and the findings were added to a shared database that helped to indicate the duplicated items.
The next step in the selection of relevant papers was the analysis of the title and the abstract. If a paper was irrelevant based on the title and/or the abstract, it was excluded (In case of doubt, the researcher read the full text). If the title and abstract were in the scope, the paper became a candidate.
If a paper remained relevant, its main properties were noted to support our findings. After the full-text reading of the candidates, we had a set of papers conforming to the inclusion criteria, and the entire procedure yielded 32 publications for a detailed analysis. At this point, the bibliometric analysis was written, and the methodology for analysis was developed.
@@ V-A Taxonomy
As for the nature of monitored data (see the “Monitored Data” column in Table III), we have used the term network traffic for studies which have proposed network monitoring methods without considering the application layer protocol in network packets. In the rest of the article, we will call such methods network traffic monitoring methods. Almost all the studies published until 2022 were not focusing on monitoring network traffic, with the work by Ptasinski et al. [58] being the only exception. Starting from 2023, a new trend can be observed where a significant proportion of papers have described such methods [39, 40, 42, 43, 35, 31, 34]. However, these studies have been limited to specific environments like maritime IoT networks. Also, according to Figure 2, the majority of papers that have proposed the analysis of application layer data have employed it for security monitoring of navigational systems, whereas the analysis of OT data is generally used in the context of CPS.
@@ V-B Security monitoring methods for distributed maritime systems and networks
Kumar et al. [42] proposed a monitoring system that uses Cyber Threat Intelligence (CTI) extracted with the DL-based method, Long Short-Term Memory (LSTM) based Variational Autoencoder. The system applied the Bi-directional Gated Recurrent Unit (Bi-GRU) method for detecting attacks addressing IoT devices. According to the authors, the proposed system featured a high performance, having a better accuracy than several other similar methods from non-maritime research literature. The main drawback of the study is the fact that no maritime-specific dataset was used for the validation of the method, but rather the TON-IoT dataset [^2] https://research.unsw.edu.au/projects/toniot-datasets which contains network traffic of consumer IoT devices. As one of the limitations, the study mentioned that the proposed DL-based method is computationally more expensive than traditional ML algorithms.
Liu et al. [39] proposed a CNN-MLP based Intrusion Detection model for MTSs (Maritime Transportation Systems) which was trained via Federated Learning (FL), also called FedBatch by the authors. In the mentioned model, a Convolutional Neural Network (CNN) was utilised for data feature extractions, while Multilayer Perception (MLP) was used for attack classification to locally detect intrusion on each vessel. The study targeted IoT-based MTSs and endeavoured to address the limitations inherent in conventional learning algorithms, which consume an amount of computing resources that can not be provided in resource-constrained IoT environments. The study aimed to protect data privacy by not sharing the training data obtained locally in vessels with the cloud. Instead, the models were induced locally in the remote devices, and their parameters were shared with the cloud. The authors proposed a robust and efficient model for MTS security monitoring by training and testing their model on the generic NSL-KDD dataset. [^3] https://www.unb.ca/cic/datasets/nsl.html
Ptasinski et. al [58] presented a Maximum Entropy Estimation based statistical ciphertext (CT) flow analysis mechanism to detect malicious flooding network attacks. The authors specifically focused on protecting (Quality of Service) QoS in the CT domain against DoS attacks in the context of the Automated Digital Network System (ADNS) INC III, a critical network system of the US Navy for unclassified and secret information exchange. The proposed approach detected anomalies in the network by using anomaly-free packets as training data and estimated their distribution as a baseline which represented normal traffic. For anomaly detection purposes, the network traffic was observed and its distribution was compared with the baseline distribution. Furthermore, the study provided an attack response mechanism by limiting network traffic. However, in some cases, normal traffic flow was also unnecessarily limited.
@@
Tiwari et al. [47] proposed a supervised ML approach to secure sensor networks against cyber attacks and adversarial activities. The work focused on maritime IoT-based systems, addressing the attack detection in the network layer of such infrastructures. The paper introduced the data collection, pre-processing, model training, method optimisation, testing and its results. According to the results, the Enhanced Light Gradient Boosting Machine (Light-GBM) technique delivered excellent results on low computational cost and network bandwidth usage, which is essential in the complex marine IoT environment, that is also limited from the communication bandwidth’s viewpoint. The experiments and the IoT attacks classification were conducted on the Distributed Smart Space Orchestration System (DS2OS) dataset, [^5] https://www.kaggle.com/datasets/francoisxa/ds2ostraffictraces an open-source but not maritime-related dataset.
@@
Abbad et al. [32] suggested the use of stochastic semi-Markov processes for monitoring NMEA messages in maritime IoT networks, so that DDoS attacks in such networks could be identified in a timely fashion. The system proposed by the authors involved data collection from IoT devices and the detection of device states which reflect the likelihood of a security incident. Also, Markov chains were used to model state transitions in the network for identifying anomalies which can indicate DDoS attacks. The proposed system was evaluated by simulating normal traffic and DDoS attacks in maritime IoT systems, and according to the authors, the system demonstrated excellent performance with a low rate of false positives and false negatives.
@@ V-C Security monitoring methods for navigational systems
Kontopoulos et al. [54] presented a distributed architecture capable of detecting spoofing and falsification attacks in streams of AIS data in real time. Average speed is measured on the shortest path between positions obtained from two consecutive AIS broadcast messages. Then, this value constitutes the basis of the spoof detection approach implemented in this study. The architecture was of the Master-Workers type, with the Master distributing the AIS messages to Workers for analysis in an efficient manner incorporating load balancing techniques. The approach was validated by generating a 43,912,236 AIS messages dataset from marinetraffic.com and performing simulations by attacking the dataset in various ways. The validation dataset has not been made public. The reduction of false positives as well as the extension to more attack types was mentioned to be part of future work.
@@
The study by Boudehenn et al. [15] focused on the detection of GPS spoofing. For that purpose, an anomaly detection approach was proposed which employed a One-Class Support Vector Machine (OC-SVM) and analysed NMEA messages that carried GPS information. The authors selected OC-SVM because it was easy to implement and did not require much computational resources, making it suitable for deployment on low-end hardware. During the experiments, the authors trained OC-SVM on data collected from a maritime simulator and also on real-life data collected on a small boat. For implementing GPS spoofing during the experiments, the authors utilised a device which generated a GPS signal carrying wrong GPS information. According to the experiments, the OC-SVM based anomaly detection method featured a high recall and precision. To demonstrate the lightweight nature of the method and its low computational cost, the authors implemented the method on a Raspberry Pi based monitoring device that was suitable for deployment on real-life ships. Such evaluation of the actual computational cost of the proposed method and its suitability for real-life deployment sets this study apart from many other studies.
Similarly to the previous paper by Boudehenn et al. [15], the study by Spravil et al. [41] focused on the detection of GPS spoofing with the analysis of relevant NMEA messages. According to the authors, the analysis of NMEA messages allowed for the development of dedicated monitoring modules which could be easily integrated into the ship’s network. To demonstrate the viability of that approach, the authors publicly released the implementation of the MANA (MAritime Nmea-based Anomaly detection) framework. The framework combined unsupervised anomaly detection methods into an ensemble, and to identify suitable methods, the authors investigated various GPS spoofing methods, with a number of methods originating from other domains than maritime (e.g., aviation and mobile phone networks). As a result of the investigation, five methods were selected for the ensemble. To assess the performance of the MANA framework, the authors created the MARSIM dataset which contained a large number of different GPS spoofing attacks and normal GPS data. Also, the dataset was released into the public domain. [^10] https://github.com/fkie-cad/mana According to the experiments, the ensemble of five methods featured a higher recall than each method (the ensemble regarded the anomaly as detected if any of the five methods provided a relevant indication). As one of the limitations of their approach, the authors identified the need for well-chosen thresholds for each method to avoid false positive alerts.
@@
The study by Longo et al. [37] described a number of attack scenarios for manipulating the maritime radar image which involved the misuse of the ASTERIX protocol. For detecting these attacks, the authors proposed a rule-based system which utilised candidate policies, where each candidate policy described the normal operating conditions of a radar. For making candidate policies adjustable for specific environments, the policies contained variables which were automatically set by the system. Also, if some candidate policies were not applicable in the given environment, they were automatically excluded from consideration. Selected policies were used for the real-time analysis of ASTERIX messages, and alerts were raised on policy violations. According to the study, the proposed system featured a high attack detection rate with a small number of false positives. In addition, the computational cost of the system was low, making it suitable for resource-constrained maritime environments. The experiments conducted in the study were supported by the MaCySTe testbed [38], which will be described in Section V-F.
To avoid data privacy issues of conventional ML and DL approaches that typically require centralising data on a single server, Hossain et al. [30] presented a FL-based approach with the aim to enhance the detection and classification of cyber attacks against marine radar systems. The method constructed a global model by utilising the collective learning ability of several distant clients, which involved local training at clients without sharing sensitive data. For evaluation, the RadarPWN dataset was used and partitioned into numerous subsets to represent different data sources in a collective learning environment. The authors provided a comprehensive evaluation of their approach, with high values reported for each performance metric. Additionally, the authors evaluated several DL-based approaches for the attack detection component of their system, choosing CNN as the best fit. The authors stated some limitations of their study, e.g., the need for reliable communication between clients and the central server, which might not always be practical in maritime operations due to connectivity issues.

### v3

# A Systematic Literature Review of Cyber Security Monitoring in Maritime

@@ Abstract
In recent years, many cyber incidents have occurred in the maritime sector, targeting the information technology (IT) and operational technology (OT) infrastructure. One of the key approaches for handling cyber incidents is cyber security monitoring, which aims at timely detection of cyber attacks with automated methods. Although several literature review papers have been published in the field of maritime cyber security, none of the previous studies has focused on cyber security monitoring. The current paper addresses this research gap and surveys the methods, algorithms, tools and architectures used for cyber security monitoring in the maritime sector. For the survey, a systematic literature review of cyber security monitoring studies is conducted following the Preferred Reporting Items for Systematic reviews and Meta-Analyses (PRISMA) protocol. The first contribution of this paper is the bibliometric analysis of related literature and the identification of the main research themes in previous works. For that purpose, the paper presents a taxonomy for existing studies which highlights the main properties of maritime cyber security monitoring research. The second contribution of this paper is an in-depth analysis of previous works and the identification of research gaps and limitations in existing literature. The gaps and limitations include several dataset and evaluation issues and a number of understudied research topics. Based on these findings, the paper outlines future research directions for cyber security monitoring in the maritime field.
@@
cyber security monitoring in maritime, cyber security monitoring, maritime cyber security, maritime, literature review
@@ I Introduction
One of the key approaches for countering cyber attacks is cyber security monitoring (or security monitoring for the sake of brevity), which aims at detecting cyber attacks with automated methods in a timely fashion, allowing to mitigate the attacks in a manual, semi-automated, or fully automated way. For example, network and host-based Intrusion Detection Systems (IDSs) are widely acknowledged security monitoring technologies for real-time surveillance of computer networks and individual hosts to detect cyber attacks. Also, most security-aware organisations operate a Security Operations Centre (SOC) for collecting data from security monitoring tools to a central Security Information and Event Management (SIEM) system where it is processed and then presented to human security analysts for further action [16]. In addition to existing industrial solutions, many experimental approaches like Machine Learning (ML) based algorithms have been proposed for cyber security monitoring in recent academic literature (see the domain overview paper [17] for more details).
@@ II Related work
The SLR paper [19] focused on cyber attacks in maritime supply chain networks (MSCN), providing recommendations on the prevention and mitigation practices. The paper introduced a comprehensive taxonomy for cyber attacks which categorised them based on a number of factors, including the hacker group, attack type, affected systems, geopolitical impact, etc. According to the study, the key cyber security threats to MCSN include malware, Distributed Denial of Service (DDoS) and Denial of Service (DoS) attacks. For prevention and mitigation, the paper proposed a number of technical and policy-level measures (e.g., the use of multi-factor authentication, standardisation of cyber security practices, etc.), and practices for increasing cyber security awareness (e.g., development of training courses about attack prevention). The study briefly mentioned some security monitoring technologies without a deeper analysis.
@@
In [22], the papers on cyber security in the maritime domain were reviewed, describing common trends in relevant literature. As commonly occurring research themes, the authors identified the impact assessment for maritime cyber threats, recognition of maritime cyber threats (e.g., through vulnerability management tools or through interviews with human personnel), and categorisation and mitigation of maritime cyber risks. The study also presented recommendations for risk mitigation, and discussed gaps in existing literature and avenues for further research. As a supplementary material to their paper, the authors published a list of 21 maritime cyber attacks that have received media coverage and public attention, providing a short description of each cyber attack. Although the study briefly mentioned couple of security monitoring approaches for cyber risk mitigation (e.g., the use of ML methods for AIS anomaly detection), these approaches were not elaborated further.
The overview paper [7] focused on maritime cyber security threats, analysing relevant literature in the field. According to the authors, existing papers do not provide enough data on past cyber security incidents and this complicates research on modelling these attacks. In addition, the authors stated that the research on maritime cyber attack data-sharing methods has received little attention so far. Furthermore, it was stated that existing studies have not focused on assessing the financial impact of cyber attacks. The authors also noted that the current educational system does not prepare maritime professionals well enough to handle cyber threats. Finally, the authors pointed out that existing governmental and international policies do not regulate maritime cyber security well enough. For example, regulations should facilitate timely reporting of cyber security incidents and cooperation between maritime companies to increase knowledge sharing on the nature and mitigation of cyber attacks.
Erbas et al. [20] performed an SLR that critically examined and compared existing threat modelling and risk assessment methods in ship cyber security, developing a taxonomy for them. The authors analysed 25 scientific papers to understand the evolving landscape of cyber security practices for crewed and autonomous ships. Significant inconsistencies were observed in current approaches and key challenges were identified, underscoring the urgent need for standardised threat modelling and risk assessment frameworks that consider also the special requirements of autonomous ships. Furthermore, the authors highlighted disparities, as well as inadequacy in tool support for the reviewed threat modelling and risk assessment methods. The need for more systematic validation of expert knowledge was also deemed critical. The topic reviewed by Erbas et al. is complementary to security monitoring, as threat modelling and risk assessment can provide direct inputs for enhanced monitoring solutions.
@@
The bibliometric review [23] analysed the research literature on the risk, reliability, and safety of autonomous ships. Whereas cyber security was not the primary focus of this paper, it was identified as an important part of the ship’s safety. Also, the study indicated that the cyber security of autonomous vessels has started to receive more attention from researchers recently. However, the paper did not include the treatment of cyber security monitoring topics.
Another study by Ben Farah et al. [4] reviewed cyber incidents in the maritime domain and provided a concise bibliographic analysis. The authors provided an extensive elaboration of maritime sub-systems located in vessels and ports, in addition to a review of the potential cyber attacks targeting these systems. Similarly to previous studies described in this section, the authors did not focus on cyber security monitoring issues.
@@ III-B Survey Method
During the development of this literature review the PRISMA protocol [25, 26] was followed. PRISMA is a widely used protocol for conducting an SLR and has been employed by several maritime cyber security literature review papers discussed in Section II [3, 20, 21, 19]. According to the PRISMA protocol, an SLR should have clearly defined objectives or research questions (see Section III-A), eligibility criteria for relevant literature (discussed in Section III-B1), a search strategy for finding relevant studies and selection process (discussed Section III-B2), and data extraction process (discussed in Section III-B3).
@@ III-B2 Search Strategy
The paper selection process has been summarised in Figure 1. The automated search from all libraries resulted in 1168 papers. After the library search, the four co-authors processed one library each in the form of independent work, and the findings were added to a shared database that helped to indicate the duplicated items. After the removal of the duplicates, 986 papers remained for further analysis.
Figure 1: The use of PRISMA protocol for identifying relevant publications
The next step in the selection of relevant papers was the analysis of the title and the abstract. If a paper was irrelevant based on the title and/or the abstract, it was excluded (In case of doubt, the researcher read the full text). If the title and abstract were in the scope, the paper became a candidate. This part of the selection resulted in 115 candidates.
If a paper remained relevant, its main properties were noted to support our findings. After the full-text reading of the candidates, we had a set of papers conforming to the inclusion criteria. The procedure ended in 29 papers and snowballing identified 3 additional studies, yielding 32 publications for a detailed analysis. At this point, the bibliometric analysis was written, and the methodology for analysis was developed.
@@ V-A Taxonomy
As for the nature of monitored data (see the “Monitored Data” column in Table III), we have used the term network traffic for studies which have proposed network monitoring methods without considering the application layer protocol in network packets. In the rest of the article, we will call such methods network traffic monitoring methods. Almost all the studies published until 2022 were not focusing on monitoring network traffic, with the work [58] being the only exception. Starting from 2023, a new trend can be observed where a significant proportion of papers have described such methods [39, 40, 42, 43, 35, 31, 34]. However, these studies have been limited to specific environments like maritime IoT networks. Also, according to Figure 3, the majority of papers that have proposed the analysis of application layer data have employed it for security monitoring of navigational systems, whereas the analysis of OT data is generally used in the context of CPS.
@@ V-B Security monitoring methods for distributed maritime systems and networks
In [42], a monitoring system was proposed that used Cyber Threat Intelligence (CTI) extracted with the DL-based method, Long Short-Term Memory (LSTM) based Variational Autoencoder. The system applied the Bi-directional Gated Recurrent Unit (Bi-GRU) method for detecting attacks addressing IoT devices. According to the authors, the proposed system featured a high performance, having a better accuracy than several other similar methods from non-maritime research literature. The main drawback of the study is the fact that no maritime-specific dataset was used for the validation of the method, but rather the TON-IoT dataset [^2] https://research.unsw.edu.au/projects/toniot-datasets which contains network traffic of consumer IoT devices. As one of the limitations, the study mentioned that the proposed DL-based method is computationally more expensive than traditional ML algorithms.
Liu et al. [39] proposed a CNN-MLP based Intrusion Detection model for MTSs which was trained via Federated Learning (FL), also called FedBatch by the authors. In the mentioned model, a Convolutional Neural Network (CNN) was utilised for data feature extractions, while Multilayer Perception (MLP) was used for attack classification to locally detect intrusion on each vessel. The study targeted IoT-based MTSs and endeavoured to address the limitations inherent in conventional learning algorithms, which consume an amount of computing resources that can not be provided in resource-constrained IoT environments. The study aimed to protect data privacy by not sharing the training data obtained locally in vessels with the cloud. Instead, the models were induced locally in the remote devices, and their parameters were shared with the cloud. The authors proposed a robust and efficient model for MTS security monitoring by training and testing their model on the generic NSL-KDD dataset. [^3] https://www.unb.ca/cic/datasets/nsl.html
In [58], a method for detecting network flooding attacks was presented which harnessed a Maximum Entropy Estimation based statistical ciphertext (CT) flow analysis mechanism. The authors specifically focused on protecting Quality of Service (QoS) in the CT domain against DoS attacks in the context of the Automated Digital Network System (ADNS) INC III, a critical network system of the US Navy for unclassified and secret information exchange. The proposed approach detected anomalies in the network by using anomaly-free packets as training data and estimated their distribution as a baseline which represented normal traffic. For anomaly detection purposes, the network traffic was observed and its distribution was compared with the baseline distribution. Furthermore, the study provided an attack response mechanism by limiting network traffic. However, in some cases, normal traffic flow was also unnecessarily limited.
@@
In [47], a supervised ML approach was proposed to secure sensor networks against cyber attacks and adversarial activities. The work focused on maritime IoT-based systems, addressing the attack detection in the network layer of such infrastructures. The paper introduced the data collection, pre-processing, model training, method optimisation, testing and its results. According to the results, the Enhanced Light Gradient Boosting Machine (Light-GBM) technique delivered excellent results on low computational cost and network bandwidth usage, which is essential in the complex marine IoT environment, that is also limited from the communication bandwidth’s viewpoint. The experiments and the IoT attacks classification were conducted on the Distributed Smart Space Orchestration System (DS2OS) dataset, [^5] https://www.kaggle.com/datasets/francoisxa/ds2ostraffictraces an open-source but not maritime-related dataset.
@@
The paper [32] suggested the use of stochastic semi-Markov processes for monitoring NMEA messages in maritime IoT networks, so that DDoS attacks in such networks could be identified in a timely fashion. The system proposed by the authors involved data collection from IoT devices and the detection of device states which reflect the likelihood of a security incident. Also, Markov chains were used to model state transitions in the network for identifying anomalies which can indicate DDoS attacks. The proposed system was evaluated by simulating normal traffic and DDoS attacks in maritime IoT systems, and according to the authors, the system demonstrated excellent performance with a low rate of false positives and false negatives.
@@ V-C Security monitoring methods for navigational systems
In [54], a distributed architecture was presented for detecting spoofing and falsification attacks in streams of AIS data in real time. Average speed was measured on the shortest path between positions obtained from two consecutive AIS broadcast messages. Then, this value constituted the basis of the spoof detection approach implemented in the study. The architecture was of the Master-Workers type, with the Master distributing the AIS messages to Workers for analysis in an efficient manner incorporating load balancing techniques. The approach was validated by generating a 43,912,236 AIS messages dataset from marinetraffic.com and performing simulations by attacking the dataset in various ways. The validation dataset has not been made public. The reduction of false positives as well as the extension to more attack types was mentioned to be part of future work.
@@
The study [15] focused on the detection of GPS spoofing. For that purpose, an anomaly detection approach was proposed which employed a One-Class Support Vector Machine (OC-SVM) and analysed NMEA messages that carried GPS information. The authors selected OC-SVM because it was easy to implement and did not require much computational resources, making it suitable for deployment on low-end hardware. During the experiments, the authors trained OC-SVM on data collected from a maritime simulator and also on real-life data collected on a small boat. For implementing GPS spoofing during the experiments, the authors utilised a device which generated a GPS signal carrying wrong GPS information. According to the experiments, the OC-SVM based anomaly detection method featured a high recall and precision. To demonstrate the lightweight nature of the method and its low computational cost, the authors implemented the method on a Raspberry Pi based monitoring device that was suitable for deployment on real-life ships. Such evaluation of the actual computational cost of the proposed method and its suitability for real-life deployment sets this study apart from many other works.
Similarly to the previous paper [15], the study by Spravil et al. [41] focused on the detection of GPS spoofing with the analysis of relevant NMEA messages. According to the authors, the analysis of NMEA messages allowed for the development of dedicated monitoring modules which could be easily integrated into the ship’s network. To demonstrate the viability of that approach, the authors publicly released the implementation of the MANA (MAritime Nmea-based Anomaly detection) framework. The framework combined unsupervised anomaly detection methods into an ensemble, and to identify suitable methods, the authors investigated various GPS spoofing methods, with a number of methods originating from other domains than maritime (e.g., aviation and mobile phone networks). As a result of the investigation, five methods were selected for the ensemble. To assess the performance of the MANA framework, the authors created the MARSIM dataset which contained a large number of different GPS spoofing attacks and normal GPS data. Also, the dataset was released into the public domain. [^10] https://github.com/fkie-cad/mana According to the experiments, the ensemble of five methods featured a higher recall than each method (the ensemble regarded the anomaly as detected if any of the five methods provided a relevant indication). As one of the limitations of their approach, the authors identified the need for well-chosen thresholds for each method to avoid false positive alerts.
@@
The study [37] described a number of attack scenarios for manipulating the maritime radar image which involved the misuse of the ASTERIX protocol. For detecting these attacks, the authors proposed a rule-based system which utilised candidate policies, where each candidate policy described the normal operating conditions of a radar. For making candidate policies adjustable for specific environments, the policies contained variables which were automatically set by the system. Also, if some candidate policies were not applicable in the given environment, they were automatically excluded from consideration. Selected policies were used for the real-time analysis of ASTERIX messages, and alerts were raised on policy violations. According to the study, the proposed system featured a high attack detection rate with a small number of false positives. In addition, the computational cost of the system was low, making it suitable for resource-constrained maritime environments. The experiments conducted in the study were supported by the MaCySTe testbed [38], which will be described in Section V-F.
To avoid data privacy issues of conventional ML and DL approaches that typically require centralising data on a single server, Hossain et al. [30] presented an FL-based approach with the aim to enhance the detection and classification of cyber attacks against marine radar systems. The method constructed a global model by utilising the collective learning ability of several distant clients, which involved local training at clients without sharing sensitive data. For evaluation, the RadarPWN dataset was used and partitioned into numerous subsets to represent different data sources in a collective learning environment. The authors provided a comprehensive evaluation of their approach, with high values reported for each performance metric. Additionally, the authors evaluated several DL-based approaches for the attack detection component of their system, choosing CNN as the best fit. The authors stated some limitations of their study, e.g., the need for reliable communication between clients and the central server, which might not always be practical in maritime operations due to connectivity issues.